OpenStudy (anonymous):

Which shows 52^2 - 48^2 being evaluated using the difference of perfect squares method?

OpenStudy (mathmate):

Hint: use the identity: \((a+b)(a-b)=a^2-b^2\)

OpenStudy (anonymous):

\(52^2-48^2=(52+48)(52-48)=\dots\)

OpenStudy (anonymous):

my bad 522 - 482 = (2704 + 2304)(2704 - 2304) = 2,003,200 522 - 482 = (52 + 48)(52 - 48) = (100)(4) = 400 522 - 482 = 2704 - 2304 = 400 522 - 482 = (52 - 48)2 = (4)2 = 16 the choices

OpenStudy (anonymous):

Don't you mean: \[52^2 - 48^2 = (2704 + 2304)(2704 - 2304) = 2,003,200\]\[52^2 - 48^2 = (52 + 48)(52 - 48) = (100)(4) = 400\]\[52^2 - 48^2 = 2704 - 2304 = 400\]\[52^2 - 48^2 = (52 - 48)2 = (4)2 = 16\] Its a bit misleading when you write it the way you did above...

OpenStudy (anonymous):

@Nitotheburito So which answer goes along with the method the guys wrote above?

OpenStudy (anonymous):

ah i just coppied and pasted it im sorry im sorry lo siento i dont really pay attention. ah I'm thinkin the second one?

OpenStudy (anonymous):

@ilfy214 come back to me :'( haha

OpenStudy (anonymous):

Oh sorry! And yes, thats correct... But make sure you understand it instead of just guessing